OpenStudy (anonymous):

are 3x+33 over x+11 and 3 equivalent? Using complete sentences explain why or why not

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 3x+33 }{ x+11 }=\frac{ 3(x+11) }{ x+11 }\] now u can see it babe

hero (hero):

That's border line inappropriate @itsmylife

OpenStudy (anonymous):

hmm how come, sir

hero (hero):

I don't know...ask @Preetha or @satellite about it

OpenStudy (anonymous):

lol aint no body got time for that ...bt thanks for reminding